# FedGraph: A Research Library and Benchmark for Federated Graph Learning

### v3

@@ 5.4 System Performance Monitoring and Resource Utilization
We monitor system performance across 10 training nodes and a central server node using Grafana, which visualizes real-time metrics collected via Prometheus. Figure 11 provides an overview of system behavior during node classification experiments. The first three plots show global test accuracy across training rounds for the Cora, Citeseer, and Pubmed datasets. FedGCN demonstrates significantly faster convergence and higher final accuracy than FedAvg across all datasets. Figure 11 (right) shows CPU usage over time, captured via Grafana. Lighter workloads, such as Cora and Citeseer, result in lower and less frequent CPU utilization spikes. In contrast, Pubmed and Ogbn-Arxiv induce higher and more sustained CPU usage due to their larger graph sizes and increased communication overhead. The pattern of usage spikes aligns with scheduled training rounds, reflecting real-time computational demands at scale.
@@ 5.5 Real-World Dataset with Realistic Client Data Distribution
Figure 12: Training Time and Test Accuracy on Ogbn-Papers100M under Different Batch Sizes (800 rounds).
We finally evaluate FedGraph at scale on the Ogbn-Papers100M dataset (over 50GB), one of the largest publicly available graph benchmarks. We use Hugging Face for dataset storage and partitioning, with the Trainer class managing local data loading for each client. The data is distributed across 195 clients  [^2] https://huggingface.co/FedGraph, with node counts assigned following a power law distribution based on country population sizes—mimicking realistic federated environments where larger clients hold more data.
@@ 6 Conclusion
Appendix A FedGraph Code Structure
@@
Backbone Models Module: gnn_models.py. It defines different backbone model classes. Generally, it is task-based (i.e., each task corresponds to one backbone model), but for some models, their variants are also included and the user could also switch the backbone model or define a new one by itself.
Utility Functions: utils_nc.py, utils_gc.py, utils_lp.py. In the current version, the utility functions for different tasks are located in separate Python modules, which is convenient for developing. In the later versions it might be better to use the shared ‘utils.py‘, and use different markers like “NC”, “GC”, and “LP” to distinguish them.
Appendix B FedGraph API and Runners
When calling the FedGraph API, as shown in Figure 13, Users can simply specify the name of the dataset and algorithm, the API will then call the corresponding data loader class to generate the required data, which is then automatically fed into the appropriate algorithm runner for tasks like Node Classification, Graph Classification, or Link Prediction. Additionally, users can seamlessly add their own datasets or federated graph learning algorithms if needed, as long as they satisfy the form requirements for the specified task.
@@ Appendix B FedGraph API and Runners
Appendix C Supported Algorithms and Datasets
@@
Ogbn-Arxiv, Ogbn-Products, Ogbn-MAG Hu et al. (2020)
@@
DD, PROTEINS, COLLAB,
IMDB-BINARY, IMDB-MULTI Xie et al. (2021)
Link Prediction FourSquare Yang et al. (2019), WyzeRule Kamani et al. (2024)
@@
Node Classification FedAvg Li et al. (2019) GCN
@@
BNS-GCN Wan et al. (2022) GCN
FedSage+ Zhang et al. (2021) GraphSage
FedGCN Yao et al. (2024) GCN,GraphSage
@@
FedAvg Li et al. (2019) GIN
FedProx Li et al. (2020) GIN
GCFL Xie et al. (2021) GIN
GCFL+ Xie et al. (2021) GIN
GCFL+dWs Xie et al. (2021) GIN
Link Prediction FedAvg Li et al. (2019) GCN
STFL Lou et al. (2021) GCN
FedGNN+ Gürler & Rekik (2022) GCN
@@
Appendix D Runner Workflow
D.1 Runner Workflow for Node Classification Task
@@
D.2 Runner Workflow for Graph Classification Task
@@
D.3 Runner Workflow for Link Prediction Task
@@
Appendix E System Measurement Dashboard
In Figure 15, the FedGraph project monitors system performance by integrating Prometheus and Grafana within a Kubernetes cluster. Prometheus collects metrics from Ray clusters, while Grafana visualizes these metrics through embedded panels, allowing real-time monitoring and analysis.
Figure 15: FedGraph Monitoring System.
@@ F.1 Parameter Configuration
In Table 4, we present the key parameters for configuring CKKS homomorphic encryption. The selection of the parameters is based on the dataset size and desired security level. Different parameter combinations create tradeoffs between computational overhead, communication cost, and precision. We present the default setting and parameters selection range to guide the user in selecting an appropriate combination that achieves the balance between objectives.
@@ F.2 Microbenchmark
We then provide the microbenchmark of HE on federated graph training in Table 5. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings in Cora. For CKKS parameters, we evaluate different polynomial modulus degrees (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds.
Dynamic Precision: We adjust encryption parameters based on graph sizes and the numerical precision needed. For graphs like Cora, a polynomial modulus degree of 16384 with precision $2^{40}$ satisfies the ideal accuracy, while increased value provides more precise security protection.
@@
When comparing among datasets, we observe that HE maintains equivalent accuracy across different parameter selections, as long as they satisfy the modulus requirement. If a smaller than required parameter size is used, the accuracy drops sharply which indicates not valid encryption.
@@
Table 5: Microbenchmark of FedGCN under Homomorphic Encryption with different CKKS scheme parameters. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings across three datasets (Cora, Citeseer, PubMed). For CKKS paramaters, we evaluate different polynomial modulus degree (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for encrypted version shows pre-train/training/total times respectively. Communication costs (Comm_cost) include both pre-training and training rounds. Plain-text FedGCN serves as the baseline for comparison. Communication cost are measured for pre-train communication and training rounds separately.
@@ Appendix G Additional Experiments
G.1 Mini-Batch for Large Dataset
FedGraph inherently supports a mini-batch graph splitter, enabling efficient graph-level processing. Our experiments on the ogbn-arxiv dataset examined the impact of varying batch sizes (16, 32, 64) on test accuracy, training time, memory usage, and communication cost across 10 clients in a federated setting.
Results in Figure 16 show test accuracy peaks at a batch size of 64, with improvements leveling off beyond 32. Training time and memory demands increase with batch size, though pretrain memory remains stable.
Figure 16: Pretrain, Train time, Communication Cost, Accuracy of FedGCN on Ogbn-Arxiv in Different Batch Sizes (200 rounds)
G.2 Increasing the Number of Clients
Figure 17: Training Time, Communication Cost, Test Accuracy on Ogbn-Arxiv in Large Number of Clients. All experiments run on 10 AWS instances. 1000 trainers take long since it runs sequentially on 10 instances.
To better test scalability and fit real-world data, we increase the number of clients to 1000. In Figure 17, we observe that as the number of clients increases, the overall training time grows significantly due to sequential running on 10 instances, added communication overhead and the need for increased synchronization among clients.

### v4

@@ 5.2 System Performance Monitoring and Resource Utilization
Refer to captionRefer to captionRefer to caption
Figure 11: Test Accuracy across Global Training Rounds (left, Cora/Citeseer/Pubmed) and Grafana Dashboard Showing CPU, Memory, and Network usage (right).
We monitor system performance across 10 training nodes and a central server node using Grafana, which visualizes real-time metrics collected via Prometheus. Figure 11 provides an overview of system behavior during node classification experiments. The first three plots show global test accuracy across training rounds for the Cora, Citeseer, and Pubmed datasets. FedGCN demonstrates significantly faster convergence and higher final accuracy than FedAvg across all datasets. Figure 11 (right) shows CPU, memory, and network usage over time, captured via Grafana. Lighter workloads, such as Cora and Citeseer, result in lower and less frequent CPU utilization spikes. In contrast, Pubmed and Ogbn-Arxiv induce higher and more sustained CPU usage due to their larger graph sizes and increased communication overhead. The pattern of usage spikes aligns with scheduled training rounds, reflecting real-time computational demands at scale.
@@ 5.3 Real-World Dataset with Realistic Client Data Distribution
Figure 12: Training Time, Test Accuracy, and Memory Usage for each client on Ogbn-Papers100M under Different Batch Sizes (800 rounds).
We finally evaluate FedGraph at scale on the Ogbn-Papers100M dataset (over 50GB), one of the largest publicly available graph benchmarks. We use Hugging Face for dataset storage and partitioning, with the Trainer class managing local data loading for each client. The data is distributed across 195 clients, with node counts assigned following a power law distribution based on country population sizes, mimicking realistic federated environments where larger clients hold more data.
@@
5.4 Benchmarking Scalability under Increasing Clients
We next evaluate the scalability of FedGraph by varying the number of clients in Table 2. As the number of clients increases, the overall communication cost grows substantially, eventually becoming the primary bottleneck. In contrast, the training time per client decreases since each client processes a smaller subgraph.
Table 2: Training and communication time (seconds) for datasets under varying client numbers.
Clients Cora CiteSeer PubMed OGBN-arXiv
Train Comm Train Comm Train Comm Train Comm
5 $1.39$ $1.69$ $1.58$ $2.78$ $2.08$ $1.55$ $127.71$ $4.48$
10 $1.36$ $2.78$ $1.79$ $6.55$ $1.77$ $2.60$ $45.82$ $5.95$
15 $1.56$ $3.99$ $2.40$ $9.58$ $1.57$ $3.84$ $21.77$ $7.77$
20 $1.49$ $4.87$ $2.07$ $13.62$ $1.83$ $4.63$ $17.89$ $9.24$
@@ 6 Conclusion
Appendix A Frequently Asked Questions
A.1 Client Selection
FedGraph supports two client selection methods that randomly or uniformly select clients at each round, as in server_class.py. Server-side algorithm components can also be added by modifying the server class.
1assert 0 < sample_ratio <= 1, "Sample ratio must be between 0 and 1"
2
3num_samples = int(self.num_of_trainers *ample_ratio)
4
5if sampling_type == "random":
6 selected_trainers_indices = random.sample(
7 range(self.num_of_trainers), num_samples
8 )
9elif sampling_type == "uniform":
10 selected_trainers_indices = [
11 (
12 i
13 + int(self.num_of_trainers * sample_ratio)
14 * current_global_epoch
15 )
16 % self.num_of_trainers
17 for i in range(num_samples)
18 ]
19
20else:
21 raise ValueError("sampling_type must be either ’random’ or ’uniform’")
A.2 Easy Integration with New Baselines
We agree that adding more FGL methods to FedGraph could make it more useful to users. Indeed, our main goal is to provide a library for benchmarking the real system performance of federated graph learning methods. Though we believe we have covered most state-of-the-art methods, we acknowledge that some methods from the literature are missing, so we also make the library easy to add new methods for researchers. Researchers can create a new training class based on an existing training method in the library (e.g., FedAvg, FedGCN).
A.3 Communication Overhead after HE with Low Rank
Homomorphic Encryption (HE) has sometimes significant drawbacks, in particular, high computing and communication loads. We include an HE implementation in the FedGraph framework as some prior work on federated graph learning has proposed HE as a way to preserve privacy, e.g.,  [20, 21, 22]. Thus, researchers in the field may wish to evaluate the effects of HE on various graph learning algorithms, or even evaluate new ways to combine HE with federated graph learning. As a benchmark platform, we do not advocate for or against such ideas; our goal is simply to facilitate the evaluation of federated graph learning algorithms that require HE implementations. For example, researchers may wish to quantitatively compare the training and communication time of their algorithms with and without HE in order to precisely measure the overhead induced by using HE. We will clarify this point in the revised manuscript.
Our work on low-rank HE is meant to serve as an example case study of the types of research that FedGraph enables with HE. We fully agree that it does not reduce the overhead of communication during training. However, FedGraph also facilitates implementing low-rank HE schemes for the aggregation process, such as FedPara [23].
A.4 Comparison with Other Framework
The FedGraphnn and FederatedScopeGNN libraries have not been maintained since 2023. When running FedAvg, due to using the same graph training library, PyG, and the distributed setup, FedGraph has a similar run time as FedGraphnn and FederatedScopeGNN. As shown in Table 1, the main advantage is supporting new algorithms, homomorphic encryption, system-level profilers, and large-scale optimization like low rank, client selection, mini-batch, etc. Given the optimization methods (e.g., low rank in Figure 7), FedGraph can run much faster than these frameworks.
A.5 Support Differential Privacy and Homomorphic Encryption
FedGraph supports differential privacy(DP) for aggregation as an option in configuration. Our implementation of DP achieves comparable performance to both the plaintext version and HE without having an accuracy loss. Table 3 provides a comparison of different matrix running FedGCN with Cora using plaintext, HE, and DP. Results are averaged over 5 runs. Both HE and DP protect the pre-training or training communication without exposing the raw data at the server in different approaches. This meets our goal of presenting FedGraph to provide user with the flexibility of using and choosing the privacy mechanism that best fits their specific needs.
Framework Pre-train Communication (MB) Pre-train Time (s) Total Time (s) Accuracy
Plaintext 56.61 4.91 12.08 0.793
HE 1208.87 17.49 40.91 0.791
DP 57.69 5.60 13.09 0.792
Table 3: Comparison of privacy preservation methods in terms of pre-train communication cost, pre-train time, total time, and accuracy.
A.6 Why HE Affects Model Accuracy
HE performance is highly sensitive to hyperparameter settings. For example, in Table 4 and Table 5, we vary the Polynomial Modulus Degree to balance security and computational capacity. A higher degree supports more complex computations but incurs greater overhead, while a lower degree restricts the allowable computation depth before noise dominates, leading to truncated or simplified operations that degrade model accuracy. In some cases, decryption itself becomes inaccurate, with the error magnitude determined by both the chosen hyperparameters and the encrypted values. Since the decrypted outputs are used for graph model training, such errors accumulate and ultimately reduce training accuracy.
A.7 Run on Edge Devices
In this paper, we focus on AWS cloud experiments as many of our envisioned use cases for federated graph learning–including learning on medical record data stored at different hospitals, or user-product consumption data stored in different countries–falls into the cross-silo federated learning paradigm, where clients are likely large servers or computers (hospitals or countries respectively, in the two examples above). FedGraph could also be run on Linux-based mobile or Internet of Things devices, e.g., Jetson Nanos, as might be appropriate for mobile applications like wearable health sensing.
Appendix B FedGraph Code Structure
@@
Backbone Models Module: gnn_models.py. It defines different backbone model classes. Generally, it is task-based (i.e., each task corresponds to one backbone model), but for some models, their variants are also included, and the user could also switch the backbone model or define a new one themselves.
Utility Functions: utils_nc.py, utils_gc.py, utils_lp.py. In the current version, the utility functions for different tasks are located in separate Python modules, which is convenient for development. In the later versions, it might be better to use the shared ‘utils.py‘, and use different markers like “NC”, “GC”, and “LP” to distinguish them.
Appendix C FedGraph API and Runners
When calling the FedGraph API, as shown in Figure 13, Users can simply specify the name of the dataset and algorithm. The API will then call the corresponding data loader class to generate the required data, which is then automatically fed into the appropriate algorithm runner for tasks like Node Classification, Graph Classification, or Link Prediction. Additionally, users can seamlessly add their own datasets or federated graph learning algorithms if needed, as long as they satisfy the form requirements for the specified task.
@@ Appendix C FedGraph API and Runners
Appendix D Supported Algorithms and Datasets
@@
Ogbn-Arxiv, Ogbn-Products, Ogbn-MAG [9]
@@
$D$, PROTEINS, COLLAB,
IMDB-BINARY, IMDB-MULTI [6]
Link Prediction FourSquare [24], WyzeRule [25]
@@
Node Classification FedAvg [18] GCN
@@
BNS-GCN [26] GCN
FedSage+ [7] GraphSage
FedGCN [8] GCN,GraphSage
@@
FedAvg [18] GIN
FedProx [27] GIN
GCFL [6] GIN
GCFL+ [6] GIN
GCFL+dWs [6] GIN
Link Prediction FedAvg [18] GCN
STFL [28] GCN
FedGNN+ [29] GCN
@@
Appendix E Runner Workflow
E.1 Runner Workflow for Node Classification Task
@@
E.2 Runner Workflow for Graph Classification Task
@@
E.3 Runner Workflow for Link Prediction Task
@@ F.1 Parameter Configuration
In Table 6, we present the key parameters for configuring CKKS homomorphic encryption. The selection of the parameters is based on the dataset size and desired security level. Different parameter combinations create tradeoffs between computational overhead, communication cost, and precision. We present the default setting and parameter selection range to guide the user in selecting an appropriate combination that achieves the balance between objectives.
@@ F.2 Microbenchmark
We then provide the microbenchmark of HE on federated graph training in Table 7. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings in Cora. For CKKS parameters, we evaluate different polynomial modulus degrees (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version show pre-train/training/total times, respectively. Communication costs (Comm_cost) include both pre-training and training rounds.
Dynamic Precision: We adjust encryption parameters based on graph sizes and the numerical precision needed. For graphs like Cora, a polynomial modulus degree of 16384 with precision $2^{40}$ satisfies the ideal accuracy, while an increased value provides more precise security protection.
@@
When comparing among datasets, we observe that HE maintains equivalent accuracy across different parameter selections, as long as they satisfy the modulus requirement. If a smaller-than-required parameter size is used, the accuracy drops sharply, which indicates invalid encryption.
@@
Table 7: Microbenchmark of FedGCN under Homomorphic Encryption with different CKKS scheme parameters. The experiments are conducted on a 2-layer FedGCN for node classification tasks, running 100 global rounds with default settings across three datasets (Cora, Citeseer, PubMed). For CKKS parameters, we evaluate different polynomial modulus degree (Poly_mod), coefficient modulus sizes (Coeff_mod), and precision levels. Time(s) for the encrypted version show pre-train/training/total times, respectively. Communication costs (Comm_cost) include both pre-training and training rounds. Plain-text FedGCN serves as the baseline for comparison. Communication costs are measured for pre-training communication and training rounds separately.
@@ Appendix G Additional Experiments
G.1 Increasing the Number of Clients with Fixed Computation Resources
Figure 15: Training Time, Communication Cost, Test Accuracy on Ogbn-Arxiv in a Large Number of Clients. All experiments run on 10 AWS instances. 1000 trainers take a long time since it runs sequentially on 10 instances.
To better test scalability and fit real-world data, we increase the number of clients to 1000. In Figure 15, we observe that as the number of clients increases, the overall training time grows significantly due to sequential running on 10 instances, added communication overhead, and the need for increased synchronization among clients.